OpenStudy (konradzuse):

find the derivative h(x) = x2 arctan(4x)

OpenStudy (konradzuse):

wee need the product rule for this one.

OpenStudy (konradzuse):

so u'v + v'u

OpenStudy (turingtest):

correct

OpenStudy (konradzuse):

2x * arctan(4x) + 4/1+4^2

OpenStudy (konradzuse):

so 2xarctan(4x) + \[\frac{1}{1+16}?\]

OpenStudy (konradzuse):

err 4/....

OpenStudy (turingtest):

you forgot the x

OpenStudy (konradzuse):

16x^2 oopsies.

OpenStudy (turingtest):

good then as you said, 4/...

OpenStudy (konradzuse):

\[\frac{4}{1+16x^2}\]

OpenStudy (turingtest):

times...?

OpenStudy (konradzuse):

*x^2

OpenStudy (turingtest):

you got it :)

OpenStudy (konradzuse):

so that would be \[\frac{4x^2}{1+16x^2}\] can't we cancel out 4x^2 and 16X^2?

OpenStudy (turingtest):

no, how would you do that? if you divide top and bottom by 4x^2 you get something ugly...

OpenStudy (konradzuse):

okies :P You cannot split it up though either right? That's only the numerator?

OpenStudy (turingtest):

you could split it up if you wanted with long division or partial fractions, but I see no point in that

OpenStudy (konradzuse):

okies...

OpenStudy (konradzuse):

so final answer is 2xtan(4x) + \[\frac{4x^2}{1+16x^2}\]

OpenStudy (turingtest):

arctan*

OpenStudy (konradzuse):

yeah that :P. I fail haha.

OpenStudy (turingtest):

it happens to us all :)

OpenStudy (konradzuse):

At first it said it was wrong, but I typed acr tan because I messed up at first typing artan and it turned art into something weird LOL :(